OpenStudy (anonymous):

Help me please! Graph the following linear system to determine the number of solutions. y equals negative three x plus six, six x plus two y equals twelve 0 1 2 infinitely many

OpenStudy (noorik.):

-3x+6 = y 6x+2y=12 you must divide the second equation by 2 to make it 3x+y = 6 then sum the two equations